OpenStudy (anonymous):

A grocery store sells pre-cut fruit. Bowls of one type of fruit are . Bowls of assorted fruit are . In one month, Benjamin paid for 12 bowls of fruit. How many assorted fruit bowls did he buy? A. 4 B. 5 C. 6 D. 7

OpenStudy (amistre64):

i think your cut and paste left a few things out that the parser couldnt recognize

OpenStudy (anonymous):

A grocery store sells pre-cut fruit. Bowls of one type of fruit are $4. Bowls of assorted fruit are $6. In one month, Benjamin paid $62 for 12 bowls of fruit. How many assorted fruit bowls did he buy?

OpenStudy (amistre64):

f + a = 12 ; total bowls f(4) + a(6) = 62 ; total paid for and how much they cost each

OpenStudy (amistre64):

this is a system of equations that can be solved for using elimination or substitution; or cramers rule :)

OpenStudy (amistre64):

i eliminate it, just simpler

OpenStudy (anonymous):

Okay so does f =x and a = y?

OpenStudy (amistre64):

f + a = 12 ; *-4 4f + 6a = 62 -4f -4a = -48 4f + 6a = 62 -------------- 2a = 14 a = 7 x and y are just names for variables; but yes. I perfer to name my variables according to the stuff im trying to determine: f = fruit bow a = assorted stuff

OpenStudy (amistre64):

7 assorteds, out of 12 total gives us 4 fruits

OpenStudy (anonymous):

Okay so the answer is 4 fruits?

OpenStudy (amistre64):

thats just information; the question is asking how many assorteds; and from this information we can determine it was 7 assorted

OpenStudy (anonymous):

Oh okay. Thanks for helping me understand this more :-)

OpenStudy (amistre64):

yw, good luck :)